OpenStudy (anonymous):

ATTENTION.>>>>>What is the area of a rectangle that is 4 x 4 1/2?

OpenStudy (zpupster):

 l*w= area of rect. 4*4.5 = ??

OpenStudy (anonymous):

18?

OpenStudy (photon336):

@MetalxAngel15 you would multiply the length times the width 4.5*4 45*4 = 18.0 yes

OpenStudy (anonymous):

thank you

OpenStudy (photon336):

no problem